OpenStudy (anonymous):

The speed of a sound wave in seawater is approximately 1500 m/s. If this wave is transmitted at 10 ms−1, what is its wavelength? Answer in units of cm

OpenStudy (rajat97):

i didn't get the question we need frequency of the wave to find the wavelength

OpenStudy (anonymous):

That's where I was having problems also but that's all the question gives

OpenStudy (rajat97):

is it 15000cm??

OpenStudy (anonymous):

I just checked and it said it was incorrect

OpenStudy (rajat97):

is it 0.66666 cm??

OpenStudy (anonymous):

did you round that in anyway?

OpenStudy (rajat97):

you can write it as 0.67cm

OpenStudy (anonymous):

its incorrect, Im not sure how to use the 10 ms^-1

OpenStudy (rajat97):

even i'm not sure about it

OpenStudy (mrnood):

The question is ambiguous. The speed of sound is given as 1500m/s - but the wave is transmitted at 10m/s. Both these are units of velocity. I suspect that you have either written the question wrong - or it is not set correctly. It would make sense if it said "this wave is transmitted at 10 s^−1" i.e. its frequency is 10 Hz (cycles per second) If v= 1500m/s and f= 10 Hz then L (wavelength)= 1500/10 = 150 m

OpenStudy (anonymous):

I am certain I have written the question correctly but I do understand the answer you've come to and it is wrong sadly, maybe the question was presented to us wrong

OpenStudy (mrnood):

I take it you have the answer - if you share it here then maybe we may be able to work out what they are aiming at. However - you know that frequency, wavelength and speed are simply related. You need to know 2 of them to calculate the other.

OpenStudy (mrnood):

The question requires the answer in cm SO - again on my assumption The waqvelength is 15000 cm

OpenStudy (anonymous):

I actually don't have the answer I just know that answers are wrong because I have 7 chances to get it right and I've used a couple so far and I believe that 15000 cm is one of them

OpenStudy (anonymous):

It's 15 cm. For some reason, it works if you change ms^-1 to kHz. So, velocity 1500 m/s, frequency 10 kHz or 10000 Hz. Do you have Fischler?

OpenStudy (anonymous):

Wow thank you for the help @mattdbrat and yes I do have Fischler